OpenStudy (anonymous):

solve b - 9.4 = -3.6 the decimals threw me off

OpenStudy (kuoministers):

put it in your calculator Boom done

OpenStudy (kuoministers):

Line the decimals up: 1.1 - 0.03 "Pad" with zeros: 1.10 - 0.03 Subtract: 1.10 - 0.03 1.07

OpenStudy (kuoministers):

http://www.mathsisfun.com/subtracting-decimals.html

OpenStudy (kuoministers):

read this then work it out.. When your done tell me what you got :)

OpenStudy (kuoministers):

hint: when you see a postive you subtact from other side.

OpenStudy (kuoministers):

b - 9.4 = -3.6 b = -3.6 + 9.4 b = 5.8

OpenStudy (kuoministers):

do you understand?

OpenStudy (kuoministers):

Well if this helped you pls click best answer Thanks!!

OpenStudy (hba):

@kira3 Hello there :) Do you understand what @Kuoministers Taught ?

OpenStudy (anonymous):

yes thanks